A 375ml bottle of antacid was analysed and found to contain 30g of Mg(OH)2 and 30g of Al(OH)3. The instructions on the bottle recommend a dose of 10ml of the antacid. Calculate the quantity of stomach acid that could be neutralized by taking the recommended dosage of antacid. Stomach acid has a pH of 0.900
1
Expert's answer
2020-04-10T13:35:53-0400

The neutraliation reactions are

2 H+ + Mg(OH)2 = Mg2+ + 2 H2O

3 H+ + Al(OH)3 = Al3+ + 3 H2O

Concentration of [H+] = 10(-pH) = 10(-0.9) = 0.126 M

The weight of Mg(OH)2 and Al(OH)3 in the dose are equal to 10/375*30 = 0.8 g,

the respective amounts are: Mg(OH)2 = 0.8/58.31 = 0.0137 mole and Al(OH)3 = 0.8/78 = 0.0103 mole.

The total amount of neutralized H+ ions is n = 2*0.0137 + 3*0.0103 = 0.0583 mole.

The volume of neutralized stomach acid is then: V = [H+]/n = 0.126 M/0.0583 mole = 2.16 L.

Answer: the volume of stomach acid that could be neutralized is 2.16 L